OpenStudy (anonymous):

Find all the solutions of tan)x_^2-tan(x)-6=0 on the interval 0,2pi possible answers x = arctan(3), x = pi - arctan(3), x = pi + arctan(2), x = pi - arctan(2) x = arctan(3), x = pi + arctan(3), x = 2pi -arctan(2), x = pi - arctan(2) x = arctan(3), x = pi + arctan(3), x = -arctan(2),x = -pi - arctan(2) The equation has no solutions on the interval [0, 2pi). I believe it is b because its the only one that fits the parameters.

OpenStudy (aum):

Is it \[(\tan(x))^2 - \tan(x) - 6 = 0\] ?

OpenStudy (anonymous):

yes thats what i meant

OpenStudy (aum):

\[(\tan(x))^2 - \tan(x) - 6 = (\tan(x)-3) * (\tan(x)+2) = 0 \\ \tan(x) = 3 \text{ or } \tan(x) = -2\]

OpenStudy (anonymous):

So b, since the others arent in the interval?

OpenStudy (aum):

Not clear which one is b.

OpenStudy (anonymous):

x = arctan(3), x = pi + arctan(3), x = 2pi -arctan(2), x = pi - arctan(2)

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

thanks

OpenStudy (aum):

yw.